could you please help me with this word problem dealing with probability?
a family owns 2 cars, a 12 yr old geo and a 1 yr old prizm. on any given morning, the probabilities that the cars will start are 0.98 and 0.84 for the geo and prism, respectfully. if the probability at least one car will start Is 0.9968, what is the probability that both will start.
thank you so much for your time and help!

Found 2 solutions by sudhanshu_kmr, edjones:
Answer by sudhanshu_kmr(1152)


probability that the geo car will start i.e P(A)= 0.98


probability that the prizm car will start i.e P(B)= 0.84

probability that the at least one car will start i.e P(A U B)= 0.9968

using .... P(AUB) = P(A) + P(B) - P(A n B)

probability that both car will start i.e P(A n B)= P(A) + P(B) - P(A U B)
= 0.98 +0.84 - 0.9968
= 0.8232


Answer by edjones(8007)

Let x=the intersection of each car starting on a Venn diagram i.e. the probability that they will both start.
.98+.84-x=.9968
-x=-.8232
x=.8232 the probability that both will start.
.
Ed